# Measuring the Discrepancy between 3D Geometric Models using Directional Distance Fields

### v1

@@ V-B2 Comparisons
The numerical results are listed in Table II, where the refinement through our DirDist further improves the accruacy of MAC significantly, while after the refinement with other distance metric, the registration accuracy even decrease dynamically. Although the RE and TE of SICP are slightly better than ours, its SR is much lower than ours, showing its limitation. The visual results are shown in Fig. 10, and when dealing with the data with noise or outliers, the details of the point clouds become indistinct, our method still register the two point clouds successfully. To evaluate the performance of different methods insightly, we also used the Registration Recall (RR) under different thresholds of RE and TE to measure the registration accuracy. As shown in Fig. 11, the superiority of our DirDist is further verified.
[…]
@@ V-C1 Implementation Details
TABLE III: Quantitative comparison of non-rigid registration on the sequences from the human motion dataset [63].
@@ V-C2 Comparisons
We compared our DirDist with CD and the P2F distance. For a fair comparison, we kept the number of sampled points in CD and the P2F distance the same as the generated reference points of our DirDist. In addition, we also compared with the state-of-the-art optimization-based non-rigid registration method called AMM [4]. We computed RMSE of the estimated and ground truth vertices of the deformed source surfaces for quantitative evaluation. The results in Table III and Fig. 12 demonstrate the significant superiority of our DirDist. Additionally, we show the registration errors of all pairs in a sequence in Fig. 13 for a more comprehensive comparison, where it can be seen that the compared methods produce much larger RMSE values for pairs with significant motion, while our DirDist always works well.
@@ V-D2 Comparisons
Following [12, 13], we employed End Point Error (EPE), Flow Estimation Accuracy (Acc) with thresholds 0.05 and 0.1 (denoted as Acc-0.05 and Acc-0.1), and Outliers as the evaluation metrics. From Table IV and Fig. 14, it can be seen that our DirDist drives much more accurate scene flows than EMD, CD, and BCD under the optimization-based framework, and our DirDist further boosts the accuracy of SOTA unsupervised learning-based methods to a significant extent, demonstrating its superiority and the importance of the distance metric in 3D data modeling.
TABLE IV: Quantitative comparisons of scene flow estimation on the Flyingthings3D dataset [64] .
@@ V-E2 Comparisons
We compared our DirDist with CD and the P2F distance. For the fairness, the number of sampled points of these two baseline methods was set the same as our DirDist. We utilized the Vertex-to-Vertex Error (V2V) between the estimated and ground truth SMPL models for quantitative evaluation. The numerical and visual results are shown in Table V and Fig. 15, respectively. Obviously, our DirDist is much better than the baseline distance metrics under the scenarios of both the whole and partial canned point clouds.
TABLE V: Quantitative comparison of parametric model estimation on the selected sequences from the CAPE dataset [65].
@@ V-F1 Point Cloud to Point Cloud
We opted for the rigid registration task to conduct an ablation study on our DirDist when both models are represented with point clouds. We assessed registration accuracy under varying hyperparameters, including the number of reference points $M$, $\beta$ in Eq. (9) that modulates the confidence score, assigning smaller weights to differences in non-overlapping areas, and the standard deviation of Gaussian noise $\sigma$ controlling the distribution of generated reference points. As shown in Table VI, with the increment of the value of $M$, registration accuracy improves slightly; however, an excessive number of reference points escalates time and memory costs, making it impractical. there is a significant decline in registration accuracy when $\beta=0$, indicating its importance and necessity. Conversely, a large $\beta$ leads to a slight drop in accuracy because all reference points exhibit similar scores, making it challenging to distinguish differences introduced by overlapping and non-overlapping areas. Finally, the value of $\sigma$ has a subtle impact on the registration accuracy.
TABLE VI: Rigid registration accuracy and calculation cost under different settings. The default setting is indicated by underlining.
Setting SR (%) RE (${}^{\circ}$) TE (cm) Time (ms) GPU (MB)
@@ V-F2 Triangle Mesh to Triangle Mesh
We utilized the non-rigid registration task to conduct an ablation study on our DirDist, where both 3D models are represented with triangle meshes. We varied the values of $M$ and $\sigma$ to examine how these two hyperparameters impact registration accuracy. The results are presented in Table VII. Evidently, employing more reference points leads to higher accuracy, but it also escalates computational costs, particularly in terms of time. Therefore, in practical applications, employing an excessive number of reference points is unnecessary. When the value of $\sigma$ is extremely small, the generated reference points are distributed very closely to the surface, and thus, fail to effectively capture the differences between the two DDFs, leading to diminished accuracy.
TABLE VII: Non-rigid registration accuracy and calculation cost under different setting. The default setting is indicated by underlining.
@@ V-F3 Triangle Mesh and Point Cloud
We opted for the parametric model estimation task to conduct an ablation study of our DirDist, employing point cloud and triangle mesh models. We varied the values of $M$, $\beta$, and $\sigma$ to examine their impact on accuracy. The results are presented in Table VIII. Evidently, employing a greater number of reference points leads to higher accuracy, but at the cost of increased computation, particularly in terms of time. Therefore, in practical applications, it is unnecessary to use an excessively large number of reference points. Considering the influence of clothing, the SMPL models and the scanned data do not perfectly overlap; consequently, overly large or small values of $\beta$ can diminish accuracy. When the value of $\sigma$ is extremely small, the reference points are positioned extremely close to the surfaces, and they cannot measure the discrepancy between the two DDFs, resulting in decreased accuracy.
TABLE VIII: Comparison of parametric model estimation accuracy under different setting of DirDist. The default setup is indicated by underlining.

### v4

@@ V-B2 Comparisons
The numerical results are listed in Table III, where the refinement through our DirDist further improves the accruacy of MAC significantly, while after the refinement with other distance metric, the registration accuracy even decrease dynamically. Although the RE and TE of SICP are slightly better than ours, its SR is much lower than ours, showing its limitation. The visual results are shown in Fig. 10, and when dealing with the data with noise or outliers, the details of the point clouds become indistinct, our method still register the two point clouds successfully. To evaluate the performance of different methods insightly, we also used the Registration Recall (RR) under different thresholds of RE and TE to measure the registration accuracy. As shown in Fig. 11, the superiority of our DirDist is further verified. Additionally, Table IV presents the computational cost per iteration during the optimization process. As shown, EMD incurs the highest computational overhead due to the necessity of constructing a comprehensive bi-directional mapping between two point sets. Similarly, ARL requires significant computational resources, as it involves calculating the intersection of lines with the two point clouds. In contrast, CD, BCD, and our DirDist exhibit comparable computational costs. However, DirDist achieves superior registration results compared to these methods, demonstrating its effectiveness and efficiency in rigid point cloud registration.
@@
TABLE IV: Running time and GPU memory costs of different distance metrics in the rigid registration task.
Method EMD CD BCD [17] ARL [3] Ours
Running Time (ms/Iter.) 559 20 21 145 22
GPU Memory (MB) 1977 1721 1730 5743 1745
@@ V-C1 Implementation Details
TABLE V: Quantitative comparison of non-rigid registration on the sequences from the human motion dataset [63].
@@ V-C2 Comparisons
We compared our DirDist with CD and the P2F distance. For a fair comparison, we kept the number of sampled points in CD and the P2F distance the same as the generated reference points of our DirDist. In addition, we also compared with the state-of-the-art optimization-based non-rigid registration method called AMM [4]. We computed RMSE of the estimated and ground truth vertices of the deformed source surfaces for quantitative evaluation. The results in Table V and Fig. 12 demonstrate the significant superiority of our DirDist. Additionally, we show the registration errors of all pairs in a sequence in Fig. 13 for a more comprehensive comparison, where it can be seen that the compared methods produce much larger RMSE values for pairs with significant motion, while our DirDist always works well. Additionally, Table VI presents the computational time and GPU memory requirements per iteration during the optimization process. Specifically, CD achieves the fastest running speed, as it only involves identifying the nearest points within the point clouds. In contrast, P2F and our proposed DirDist require more computational time due to the added complexity of locating the closest point on the triangle mesh, which is inherently more complex than finding the nearest points in a point cloud.
@@
TABLE VI: Running time and GPU memory costs of different distance metrics in the non-rigid registration task.
Method CD P2F Ours
Running Time (ms/Iter.) 42 110 111
GPU Memory (MB) 3245 3264 3265
@@ V-D2 Comparisons
Following [12, 13], we employed End Point Error (EPE), Flow Estimation Accuracy (Acc) with thresholds 0.05 and 0.1 (denoted as Acc-0.05 and Acc-0.1), and Outliers as the evaluation metrics. From Table VII and Fig. 14, it can be seen that our DirDist drives much more accurate scene flows than EMD, CD, and BCD under the optimization-based framework, and our DirDist further boosts the accuracy of SOTA unsupervised learning-based methods to a significant extent, demonstrating its superiority and the importance of the distance metric in 3D data modeling. Table VIII illustrates the computational cost of various distance metrics per iteration during the optimization process. Specifically, EMD demands significantly more computing resources due to the necessity of constructing a bi-directional mapping between two point sets. In contrast, CD, BCD, and our proposed DirDist exhibit comparable computational requirements. However, our DirDist outperforms the others in terms of flow estimation accuracy, highlighting its superior effectiveness and efficiency.
TABLE VII: Quantitative comparisons of scene flow estimation on the Flyingthings3D dataset [64] .
@@
TABLE VIII: Running time and GPU memory costs of different distance metrics under the scene flow estimation task.
Method EMD CD BCD [17] Ours
Running Time (ms/Iter.) 1021 13 14 15
GPU Memory (MB) 2257 1745 1755 1767
@@ V-E2 Comparisons
We compared our DirDist with CD and the P2F distance. For the fairness, the number of sampled points of these two baseline methods was set the same as our DirDist. We utilized the Vertex-to-Vertex Error (V2V) between the estimated and ground truth SMPL models for quantitative evaluation. The numerical and visual results are shown in Table IX and Fig. 15, respectively. Obviously, our DirDist is much better than the baseline distance metrics under the scenarios of both the whole and partial canned point clouds. Additionally, Table X presents the computational resource requirements of different distance metrics during the optimization process, showing that our DirDist achieves comparable running time and GPU memory usage. This further highlights the efficiency of our proposed method in practical applications.
TABLE IX: Quantitative comparison of parametric model estimation on the selected sequences from the CAPE dataset [65].
@@
TABLE X: Running time and GPU memory costs of different distance metrics in the parametric model estimation task.
Method CD P2F Ours
Running Time (ms/Iter.) 60 64 65
GPU Memory (MB) 1813 1810 1809
@@ V-F1 Point Cloud to Point Cloud
We opted for the rigid registration task to conduct an ablation study on our DirDist when both models are represented with point clouds. We assessed registration accuracy under varying hyperparameters, including the number of reference points $M$, $\beta$ in Eq. (9) that modulates the confidence score, assigning smaller weights to differences in non-overlapping areas, and the standard deviation of Gaussian noise $\sigma$ controlling the distribution of generated reference points. As shown in Table XI, with the increment of the value of $M$, registration accuracy improves slightly; however, an excessive number of reference points escalates time and memory costs, making it impractical. there is a significant decline in registration accuracy when $\beta=0$, indicating its importance and necessity. Conversely, a large $\beta$ leads to a slight drop in accuracy because all reference points exhibit similar scores, making it challenging to distinguish differences introduced by overlapping and non-overlapping areas. Finally, the value of $\sigma$ has a subtle impact on the registration accuracy.
TABLE XI: Rigid registration accuracy and calculation cost under different settings. The default setting is indicated by underlining.
Setting SR (%) RE () TE (cm) Time (ms) GPU (MB)
@@ V-F2 Triangle Mesh to Triangle Mesh
We utilized the non-rigid registration task to conduct an ablation study on our DirDist, where both 3D models are represented with triangle meshes. We varied the values of $M$ and $\sigma$ to examine how these two hyperparameters impact registration accuracy. The results are presented in Table XII. Evidently, employing more reference points leads to higher accuracy, but it also escalates computational costs, particularly in terms of time. Therefore, in practical applications, employing an excessive number of reference points is unnecessary. When the value of $\sigma$ is extremely small, the generated reference points are distributed very closely to the surface, and thus, fail to effectively capture the differences between the two DDFs, leading to diminished accuracy.
TABLE XII: Non-rigid registration accuracy and calculation cost under different setting. The default setting is indicated by underlining.
@@ V-F3 Triangle Mesh and Point Cloud
We opted for the parametric model estimation task to conduct an ablation study of our DirDist, employing point cloud and triangle mesh models. We varied the values of $M$, $\beta$, and $\sigma$ to examine their impact on accuracy. The results are presented in Table XIII. Evidently, employing a greater number of reference points leads to higher accuracy, but at the cost of increased computation, particularly in terms of time. Therefore, in practical applications, it is unnecessary to use an excessively large number of reference points. Considering the influence of clothing, the SMPL models and the scanned data do not perfectly overlap; consequently, overly large or small values of $\beta$ can diminish accuracy. When the value of $\sigma$ is extremely small, the reference points are positioned extremely close to the surfaces, and they cannot measure the discrepancy between the two DDFs, resulting in decreased accuracy.
TABLE XIII: Comparison of parametric model estimation accuracy under different settings of DirDist. The default setup is indicated by underlining.